# Scalable Influence and Fact Tracing for Large Language Model Pretraining

### v1

@@ Abstract
Training data attribution (TDA) methods aim to attribute model outputs back to specific training examples, and the application of these methods to large language model (LLM) outputs could significantly advance model transparency and data curation. However, it has been challenging to date to apply these methods to the full scale of LLM pretraining. In this paper, we refine existing gradient-based methods to work effectively at scale, allowing us to retrieve influential examples for an 8B-parameter language model from a pretraining corpus of over 160B tokens with no need for subsampling or pre-filtering. Our method combines several techniques, including optimizer state correction, a task-specific Hessian approximation, and normalized encodings, which we find to be critical for performance at scale. In quantitative evaluations on a fact tracing task, our method performs best at identifying examples that influence model predictions, but classical, model-agnostic retrieval methods such as BM25 still perform better at finding passages which explicitly contain relevant facts. These results demonstrate a misalignment between factual attribution and causal influence. With increasing model size and training tokens, we find that influence more closely aligns with attribution. Finally, we examine different types of examples identified as influential by our method, finding that while many directly entail a particular fact, others support the same output by reinforcing priors on relation types, common entities, and names.
@@ 1 Introduction
However, the steep computational cost of applying TDA approaches to LLMs has limited work in this area. Studies applying TDA to LLM fine-tuning have achieved promising results identifying examples that influence model behavior (Akyurek et al., 2022; Park et al., 2023; Xia et al., 2024), but there is significant evidence that much of an LLM’s knowledge and capabilities originates from pretraining (Hoffmann et al., 2022; Chang & Bergen, 2024). Previous work applying TDA to pretraining has either focused on small models (Engstrom et al., 2024) or extremely few target queries, e.g. retrieving influential examples from a significantly subsampled corpus for less than 30 model predictions (Grosse et al., 2023; Choe et al., 2024). In our work, we scale TDA experiments to LLMs up to 8B parameters, thousands of model predictions, and corpora up to 160B tokens.
@@
Specifically we propose TrackStar, a gradient-based influence method that combines innovations from previous work and scales to large setups (§3), while still supporting efficient retrieval of influential examples and quantitative evaluation. At scale, our method significantly outperforms previous influence methods at retrieving pretraining examples that entail a fact (attribution) and examples that influence specific factual predictions (influence) (§5, §6). Importantly, we demonstrate a misalignment between attribution and influence; classical methods such as BM25 are better at retrieving examples that entail factual predictions, but those examples are not necessarily those that most affect model predictions (§5.1). We show that influence grows closer to attribution as models scale, both in parameters and training tokens (§5.3). To provide insights into where attribution and influence misalign, we include analyses of examples that have high influence for factual predictions in LLMs (§7). For example, rather than containing a full statement of a fact of interest, many examples appear to support priors on relation types, common entities, or names.
@@ 2 Background: Training Data Attribution
However, when applying influence methods to LLM pretraining, an additional bottleneck is the scale of model parameters and pretraining data. Given $|\theta|$ model parameters (on the order of billions), we have loss gradients $\nabla L(z)\in\mathbb{R}^{|\theta|}$ and inverse Hessian $H^{-1}\in\mathbb{R}^{|\theta|\times|\theta|}$. To rank pretraining examples, the gradient dot product must be computed between every pretraining example $z_{m}$ (on the order of millions to billions) and the target query $z_{q}$. For this reason, previous work has focused on identifying influential examples during fine-tuning (Akyurek et al., 2022; Park et al., 2023; Xia et al., 2024) or during pretraining for smaller models (Engstrom et al., 2024). Closest to our work has been that of Grosse et al. (2023) and Choe et al. (2024) which looks at billion-parameter LLMs; however, these perform retrieval on a small subset of the pretraining corpus and report only qualitative evaluation on a small number of queries.
@@ 4 Models, Datasets, and Metrics
We apply TrackStar to identify proponent examples for factual predictions in LLMs up to 8B parameters. Unless otherwise specified, we first build an index of the projected gradients for all candidate examples of interest (e.g. the pretraining set), then compute the Hessian approximation $R$ using a single pass over the data. For gradient-based methods, we perform exact scoring between each query and all candidate examples, [^2] For online retrieval, approximate nearest-neighbor methods can also be used. with no need for lexical pre-filtering (c.f. Akyurek et al., 2022; Park et al., 2023; Grosse et al., 2023). All elements of this approach have compute cost linear in model and dataset size (§2).
@@ 4.1 Models
We pretrain a decoder-only language model on two epochs of English C4 (Raffel et al., 2020) for three model sizes: 154M, 1B, and 8B parameters, using the architecture described in Chowdhery et al. (2023) and implemented using T5X (Roberts et al., 2022). For all model sizes, we use a SentencePiece tokenizer (Kudo & Richardson, 2018) trained on C4 data with vocabulary size 32K. Using this tokenizer, English C4 consists of 160B tokens across 365M examples (short passages and documents). We pretrain with batch size 1024 and sequence length 2048 for two epochs (187K steps). All model sizes are trained on the same shuffle of the pretraining data. Exact hyperparameters are in Table A.2. The 154M, 1B, and 8B models reach eval losses (log-perplexities) of $2.34$, $1.99$, and $1.77$ respectively. We focus primarily on the 8B model, but we study the smaller models for dimensionality ablations (§5.2) and scaling (§5.3).
@@ 4.2 Fact Tracing Dataset
We focus on identifying proponent examples for factual predictions in LLMs. For factual recall prompts, we use the filtered T-REx dataset from KILT (Petroni et al., 2021), which consists of entity-relation-entity triples such as (Carleton College, country, USA). We merge the KILT dataset back into the original T-REx dataset (Elsahar et al., 2018), obtaining a dataset of facts with corresponding entity IDs, surface form aliases, and Wikipedia abstract sentences containing each fact. [^3] Because T-REx is automatically scraped, there is some inherent noise in the “ground truth” labels. For each fact, there are an average of 2.5 “ground-truth” sentences out of 19.6M sentences total. After filtering out ambiguous prompts (e.g. multiple correct target entity IDs), we have a total of 1.2M fact triples covering 97 relation types. Following Kandpal et al. (2023), we also annotate facts with the count (fact frequency) of pretraining examples from C4 that mention both entities.
We manually write natural language templates for all 97 relation types, so that a left-to-right LLM can predict the target fact as a completion, e.g. “Carleton College is located in the following country: USA”. We mark predictions as correct using string matching after lowercasing and stopword removal, considering multiple possible correct surface form aliases as provided by KILT. Random chance accuracy on this task is close to 0% due to the large number of possible completions; our 154M, 1B, and 8B models achieve $16.3\%$, $28.0\%$, and $32.4\%$, respectively. For factual queries in all following evaluations, we use a fixed sample of 5415 facts drawn randomly from this set but balanced for fact frequency and down-sampling for common target entities (such as “USA” or “English”), and limiting facts that are incorrectly predicted by all models. Each query contains an input prompt and ground truth target text. Dataset details and pre-processing are described in §A.2.
@@ 4.3 Evaluation Metrics
Attribution metrics: MRR and Recall@10. These measure whether we retrieve examples that logically support (entail) a fact. For T-REx evaluations (§5), entailing Wikipedia abstract sentences are annotated in the dataset (§4.2). For C4 evaluations (§6), because passages are not annotated, we use an entailment model to score whether a candidate passage contains factual information supporting the query. Specifically, we use a fine-tuned T5 11B (Raffel et al., 2020) model trained on ANLI (Nie et al., 2020) and synthetic data as described in Gekhman et al. (2023). [^4] This model achieves 91.5 F1 on the binary version of the MNLI validation set. We input the passage as the premise and the query as the hypothesis, and we mark a proponent as entailing when the entailment score $\geq$ 0.50. For MRR, we compute the mean reciprocal rank of the top-ranked “entailing” proponent for each fact. For recall@10, we compute the proportion of facts for which an entailing proponent appears in the top 10 proponent retrievals.
Influence metric: incremental training probability increase (tail-patch). Both of the previous metrics assume that the top-ranked proponents should be those that entail the target fact. However, these proponents are not necessarily the examples that would most influence the model to make that prediction. Following Koh & Liang (2017), most work on influence methods derives metrics from leave-one-out or other ablate-and-retrain experiments such as linear datamodeling score (Ilyas et al., 2022). However, this is not tractable for full-scale LLM pretraining, so we instead estimate the additive contribution from incremental training. Starting from the final model checkpoint and maintaining all pretraining hyperparameters, we take a single training step—which we call a tail-patch step—on a single retrieved proponent and measure the change in probability of the target sequence. We then average this change across the top 10 proponents for each query, and across all queries in the evaluation set.
@@ 5.2 Ablation Experiments
Projection dimensionality: Higher gradient projection dimensionality $d$ results in higher fidelity representations, but both the memory required to store projected gradients and the retrieval cost to compute dot products increase (linearly) with $d$. To balance these considerations, in Figure 2 (left, center), we plot how attribution (MRR) and influence (tail-patch) scores improve with projection dimensionality for TrackStar. [^9] Park et al. (2023) find that performance may decrease if projection dimensionality is too large, but our experiments do not appear to be close to that limit, likely due to the high dimensionality $|\theta|$ of LLM gradients. For smaller models (154M and 1B parameters), scores have begun to plateau around $d=2^{16}$. Although the 8B-parameter model has not yet plateaued, we restrict our experiments to $d=2^{16}$ due to memory limitations, as the index of 365M C4 examples at $d=2^{16}$ is already 87TB in size (§6). Furthermore, we note that our factual knowledge task likely requires a large number of dimensions due to the inherent difficulty of compressing individual facts about a large number of entities; fewer dimensions may be sufficient for attribution and influence for other tasks.
@@ 5.3 Influence Approaches Attribution
In §5.1, we demonstrated a distinction between attribution (identifying proponents that entail a fact, quantified by MRR and recall) and influence (identifying proponents that influence the model prediction, quantified by tail-patch scores). Interestingly, we find that TrackStar attribution scores improve as models increase in both parameters and training tokens (Figure 2, right). This indicates that as models improve, the influential examples that TrackStar identifies align more with attribution, suggesting that more capable models rely more on examples that actually entail facts for factual predictions. Of course, it does not appear that these measures converge entirely; recall from §5.1 that tail-patching on ground truth entailing proponents results in much smaller target probability changes than TrackStar proponents, even for our largest model. Thus, it appears that as models improve, they are more likely to use actual entailing examples to learn facts, but there are still many proponents that have large effects on the model despite non-entailment. We present a deeper analysis of these types of proponents in §7.
@@ 6 C4 Open Set Evaluation
BM25 0.641 0.889 +0.83%
Gecko 0.601 0.808 +0.54%
Gradient dot product +0.04%
Gradient cosine 0.261 0.451 +1.95%
TrackStar 0.346 0.581 +2.11%
Table 2: Results retrieving proponents from all of C4 (§6). Gradient cosine ablates the Hessian approximation $R$ from TrackStar, and Gradient dot product further ablates unit normalization. We omit MRR and recall from the gradient dot product results because that method almost exclusively retrieves long examples (high magnitude gradients) that are mistakenly classified as entailment by our AIS model. For that row, MRR and recall appear very close to zero. These AIS false positives are relatively uncommon for retrievals using the other methods.
@@ 6.1 C4 Open Set Results
Also in line with §5, BM25 and Gecko perform much better than TrackStar for attribution (MRR 0.641 and 0.601 vs. 0.346). Still, we note overall high MRR and recall scores given the difficulty of the task: for 58.1% of facts, TrackStar retrieves an entailing example in the top 10 out of 365M C4 examples. Additionally, TrackStar performs over 2.5$\times$ better than BM25 and Gecko at retrieving examples that influence model predictions (tail-patch score $+2.11\%$ vs. $+0.83\%$ and $+0.54\%$). This reinforces the distinction between attribution and influence: examples that entail facts—such as proponents retrieved by BM25—are often not the examples that most affect a model’s predictions.
@@ 7 Headroom Analysis
In §5 and §6 we showed that TrackStar makes significant progress towards attribution for LLM pretraining, outperforming other methods at retrieving influential pretraining examples, but it still falls short of classical baselines on the fact tracing task (attribution). To better understand why, we look in more detail at the types of proponents retrieved by our method. We find that much of the headroom can be explained by proponents which reflect priors or partial matches, multi-hop reasoning, or idiosyncrasies of the fact tracing task such as ambiguous entities or alternative correct answers. Examples of these proponents are included in Table 3 and Table 4.
@@ A.1.3 Task-Specific Hessian Approximation
We select $\lambda$ such that roughly the top 1000 out of $655362^{16}=$ 65536 task-specific components are downweighted. Concretely, we select $\lambda$ such that ${}^{\textrm{th}}$ largest singular values of $\lambda R_{\textrm{eval}}$ and $(1-\lambda)R_{\textrm{train}}$ are roughly equal. Put another way, we scale $\lambda$ such that the (monotonic) singular value curves of $\lambda R_{\textrm{eval}}$ and $(1-\lambda)R_{\textrm{train}}$ intersect at ${}^{\textrm{th}}$ component. We note that this requires larger $\lambda$ when $R_{\textrm{train}}$ is computed over C4 rather than T-REx, because C4 examples have overall larger gradients (and thus larger $R_{\textrm{train}}$ singular values) due to longer sequence lengths. We set $\lambda=0.90$ for the T-REx experiments in §5 and $\lambda=0.99$ for the C4 experiments in §6. We also find that these values work well empirically, selecting from [0.50, 0.90, 0.99, 0.999]. Determining the optimal $\lambda$ for a given task (i.e. the number of task-specific components to downweight) is an interesting direction for future investigation.

### v2

@@ Abstract
Training data attribution (TDA) methods aim to attribute model outputs back to specific training examples, and the application of these methods to large language model (LLM) outputs could significantly advance model transparency and data curation. However, it has been challenging to date to apply these methods to the full scale of LLM pretraining. In this paper, we refine existing gradient-based methods to work effectively at scale, allowing us to retrieve influential examples for an 8B-parameter language model from a pretraining corpus of over 160B tokens with no need for subsampling or pre-filtering. Our method combines several techniques, including optimizer state correction, a task-specific Hessian approximation, and normalized encodings, which we find to be critical for performance at scale. In quantitative evaluations on a fact tracing task, our method performs best at identifying examples that influence model predictions, but classical, model-agnostic retrieval methods such as BM25 still perform better at finding passages which explicitly contain relevant facts. These results demonstrate a misalignment between factual attribution and causal influence. With increasing model size and training tokens, we find that influence more closely aligns with factual attribution. Finally, we examine different types of examples identified as influential by our method, finding that while many directly entail a particular fact, others support the same output by reinforcing priors on relation types, common entities, and names. We release our prompt set and model outputs, along with a web-based visualization tool to explore influential examples for factual predictions, commonsense reasoning, arithmetic, and open-ended generation for an 8B-parameter LLM. [^1] https://github.com/pair-code/pretraining-tda
@@ 1 Introduction
However, the steep computational cost of applying TDA approaches to LLMs has limited work in this area. TDA approaches have achieved promising results identifying examples that influence LLM behavior during fine-tuning (Akyurek et al., 2022; Park et al., 2023; Xia et al., 2024), but there is significant evidence that much of an LLM’s knowledge and capabilities originates from pretraining (Hoffmann et al., 2022; Chang & Bergen, 2024). Previous work applying TDA to pretraining has either focused on small models (Engstrom et al., 2024) or extremely few target queries, e.g. retrieving influential examples from a significantly subsampled corpus for less than 100 model predictions (Grosse et al., 2023; Choe et al., 2024; Ruis et al., 2024). In our work, we scale TDA experiments to LLMs up to 8B parameters, thousands of model predictions, and corpora up to 160B tokens.
@@
Specifically we propose TrackStar, a gradient-based influence method that combines innovations from previous work and scales to large setups (§3), while still supporting efficient retrieval of influential examples and quantitative evaluation. At scale, our method significantly outperforms previous influence methods at retrieving pretraining examples that entail a fact (attribution) and examples that influence specific factual predictions (influence) (§5, §6). Importantly, we demonstrate a misalignment between attribution and influence; classical methods such as BM25 are better at retrieving examples that entail factual predictions, but those examples are not necessarily those that most affect model predictions (§5.1). We show that influence grows closer to attribution as models scale, both in parameters and training tokens (§5.3).
To provide insights into where attribution and influence misalign, we include analyses of examples that have high influence for factual predictions in LLMs (§7). For example, rather than containing a full statement of a fact of interest, many examples appear to support priors on relation types, common entities, or names.
@@ 2 Background: Training Data Attribution
However, when applying influence methods to LLM pretraining, an additional bottleneck is the scale of model parameters and pretraining data. Given $|\theta|$ model parameters (on the order of billions), we have loss gradients $\nabla L(z)\in\mathbb{R}^{|\theta|}$ and inverse Hessian $H^{-1}\in\mathbb{R}^{|\theta|\times|\theta|}$. To rank pretraining examples, the gradient dot product must be computed between every pretraining example $z_{m}$ (on the order of millions to billions) and the target query $z_{q}$. For this reason, previous work has focused on identifying influential examples during fine-tuning (Akyurek et al., 2022; Park et al., 2023; Xia et al., 2024) or during pretraining for smaller models (Engstrom et al., 2024). Closest to our work has been that of Grosse et al. (2023), Choe et al. (2024), and Ruis et al. (2024), who look at LLMs with billions of parameters; however, they perform retrieval on a small subset of the pretraining corpus and report evaluation on only a small number of queries.
@@ 4 Models, Datasets, and Metrics
We apply TrackStar to identify proponent examples for factual predictions in LLMs up to 8B parameters. Unless otherwise specified, we first build an index of the projected gradients for all candidate examples of interest (e.g. the pretraining set), then compute the Hessian approximation $R$ using a single pass over the data. For gradient-based methods, we perform exact scoring between each query and all candidate examples, with no need for lexical pre-filtering (c.f. Akyurek et al., 2022; Park et al., 2023; Grosse et al., 2023). All elements of this approach have compute cost linear in model and dataset size (§2).
@@ 4.1 Models
We pretrain a decoder-only language model on two epochs of English C4 (Raffel et al., 2020) for three model sizes: 154M, 1B, and 8B parameters, using the architecture described in Chowdhery et al. (2023). Using our tokenizer, English C4 consists of 160B tokens across 365M examples (short passages and documents). Model details are in §A.2. We focus primarily on the 8B model, but we study the smaller models for dimensionality ablations (§5.2) and scaling (§5.3).
@@ 4.2 Fact Tracing Dataset
We focus on identifying proponent examples for factual predictions in LLMs. For factual recall prompts, we use the filtered T-REx dataset from KILT (Petroni et al., 2021), which consists of entity-relation-entity triples such as (Carleton College, country, USA). We merge the KILT dataset back into the original T-REx dataset (Elsahar et al., 2018), obtaining a dataset of facts with corresponding entity IDs, surface form aliases, and Wikipedia abstract sentences containing each fact. [^3] Because T-REx is automatically scraped, there is some inherent noise in the “ground truth” labels. For each fact, there are an average of 2.5 “ground-truth” entailing sentences out of 19.6M sentences total. After filtering out ambiguous prompts (e.g. multiple correct target entity IDs), we have a total of 1.2M fact triples covering 97 relation types. Following Kandpal et al. (2023), we also annotate facts with the count (fact frequency) of pretraining examples from C4 that mention both entities.
We manually write natural language templates for all 97 relation types, so that a left-to-right LLM can predict the target fact as a completion, e.g. “Carleton College is located in the following country: USA”. We mark predictions as correct using string matching after lowercasing and stopword removal, considering multiple possible correct surface form aliases as provided by KILT. Random chance accuracy on this task is close to 0% due to the large number of possible completions; our 154M, 1B, and 8B models achieve $16.3\%$, $28.0\%$, and $32.4\%$, respectively. For factual queries in all following evaluations, we use a fixed sample of 5415 facts drawn randomly from this set but balanced for fact frequency and down-sampling common target entities (such as “USA” or “English”), and limiting facts that are incorrectly predicted by all models. Each query contains an input prompt and ground truth target text. Dataset details and pre-processing are described in §A.3.
@@ 4.3 Evaluation Metrics
Attribution metrics: MRR and Recall@10. These measure whether we retrieve examples that logically support (entail) a fact. For T-REx evaluations (§5), entailing Wikipedia abstract sentences are annotated in the dataset (§4.2). For C4 evaluations (§6), because passages are not annotated, we use an entailment model to score whether a candidate passage contains factual information supporting the query. Specifically, we use a fine-tuned T5 11B (Raffel et al., 2020) model trained on ANLI (Nie et al., 2020) and synthetic data as described in Gekhman et al. (2023). Because C4 passages can be up to 2048 tokens, we split the input C4 passages by sentences using regex matching, and we take the maximum entailment score using a sliding window of three sentences as the input premise and the factual query as the hypothesis. We mark a proponent as entailing when this entailment score $\geq$ 0.50.444Based on blinded manual annotation of 100 C4 passages marked as entailing and non-entailing respectively, we find a false positive rate of 13% and a false negative rate of 11%. Of the incorrect annotations, most (54%) are fuzzy entailments where a passage implies but technically may not entail a fact. For MRR, we compute the mean reciprocal rank of the top-ranked “entailing” proponent for each fact. For recall@10, we compute the proportion of facts for which an entailing proponent appears in the top 10 proponent retrievals.
Influence metric: incremental training probability increase (tail-patch). Both of the previous metrics assume that the top-ranked proponents should be those that entail the target fact. However, these proponents are not necessarily the examples that would most influence the model to make that prediction. Following Koh & Liang (2017), most work on influence methods derives metrics from leave-one-out or other ablate-and-retrain experiments such as linear datamodeling score (Ilyas et al., 2022). However, this is not tractable for full-scale LLM pretraining, so we instead estimate the additive contribution from incremental training. Starting from the final model checkpoint and maintaining all pretraining hyperparameters, we take a single training step—which we call a tail-patch step—on a single retrieved proponent and measure the change in probability of the target sequence. We then average this change across the top $k=10$ proponents for each query, and across all queries in the evaluation set. Results for different $k$ are reported in §A.4.2.
@@ 5.2 Ablation Experiments
Projection dimensionality: Higher gradient projection dimensionality $d$ results in higher fidelity representations, but both the memory required to store projected gradients and the retrieval cost to compute dot products increase (linearly) with $d$. To balance these considerations, in Figure 2 (left, center), we plot how attribution (MRR) and influence (tail-patch) scores improve with projection dimensionality for TrackStar. [^9] Park et al. (2023) find that performance may decrease if projection dimensionality is too large, but our experiments do not appear to be close to that limit, likely due to the high dimensionality $|\theta|$ of LLM gradients. For smaller models (154M and 1B parameters), scores have begun to plateau around $d=2^{16}$. Although the 8B-parameter model has not yet plateaued, we restrict our experiments to $d=2^{16}$ due to memory limitations, as the index of 365M C4 examples at $d=2^{16}$ is already 87TB in size (§6). Furthermore, we note that our factual knowledge task likely requires a large number of dimensions due to the inherent difficulty of compressing individual facts about a large number of entities; fewer dimensions may be sufficient for other tasks.
@@ 5.3 Influence Approaches Attribution
In §5.1, we demonstrated a distinction between attribution (identifying proponents that entail a fact, quantified by MRR and recall) and influence (identifying proponents that influence the model prediction, quantified by tail-patch scores). We find that TrackStar attribution scores improve as models increase in both parameters and training tokens (Figure 2, right). This indicates that as models improve, the influential examples that TrackStar identifies align more with attribution, suggesting that more capable models rely more on examples that actually entail facts for factual predictions. Of course, it does not appear that these measures converge entirely; even for our largest model, tail-patching on ground truth entailing proponents results in much smaller target probability changes than TrackStar proponents (§5.1). Thus it appears that as models improve, they are more likely to use entailing examples to learn facts, but there are still many proponents that have large effects on the model despite non-entailment. We present a deeper analysis of these types of proponents in §7.
@@ 6 C4 Open Set Evaluation
BM25 0.687 0.845 +0.83%
Gecko 0.636 0.826 +0.54%
Gradient dot product 0.003 0.015 +0.04%
Gradient cosine 0.252 0.393 +1.95%
TrackStar 0.338 0.515 +2.11%
Table 2: Results retrieving proponents from all of C4 (§6). Gradient cosine ablates the Hessian approximation $R$ from TrackStar, and Gradient dot product further ablates unit normalization.
@@ 6.1 C4 Open Set Results
Also in line with §5, BM25 and Gecko perform much better than TrackStar for attribution (MRR 0.687 and 0.636 vs. 0.338). Still, we note overall high MRR and recall scores given the difficulty of the task: for 51.5% of facts, TrackStar retrieves an entailing example in the top 10 out of 365M C4 examples. Additionally, TrackStar performs over 2.5$\times$ better than BM25 and Gecko at retrieving examples that influence model predictions (tail-patch score $+2.11\%$ vs. $+0.83\%$ and $+0.54\%$). This reinforces the distinction between attribution and influence: examples that entail facts—such as proponents retrieved by BM25—are often not the examples that most affect a model’s predictions.
@@ 7 Headroom Analysis
In §5 and §6 we showed that TrackStar makes significant progress towards attribution for LLM pretraining, outperforming other methods at retrieving influential pretraining examples, but it still falls short of classical baselines on the fact tracing task (attribution). To better understand why, we look in more detail at the types of proponents retrieved by our method. We find that much of the headroom can be explained by proponents which reflect priors or partial matches, multi-hop reasoning, or idiosyncrasies of the fact tracing task such as ambiguous entities or alternative correct answers. Examples of these proponents are included in Table 3 and Table 4. The full set of proponents can be viewed at https://github.com/pair-code/pretraining-tda, including proponents for additional evaluation tasks as described in §A.5.
@@ A.1.3 Task-Specific Hessian Approximation
We select $\lambda$ such that roughly the top 1000 out of $655362^{16}=$ 65536 task-specific components are downweighted. Concretely, we select $\lambda$ such that ${}^{\textrm{th}}$ largest singular values of $\lambda R_{\textrm{eval}}$ and $(1-\lambda)R_{\textrm{train}}$ are roughly equal. Put another way, we scale $\lambda$ such that the (monotonic) singular value curves of $\lambda R_{\textrm{eval}}$ and $(1-\lambda)R_{\textrm{train}}$ intersect at ${}^{\textrm{th}}$ component. We note that this requires larger $\lambda$ when $R_{\textrm{train}}$ is computed over C4 rather than T-REx, because C4 examples have overall larger gradients (and thus larger $R_{\textrm{train}}$ singular values) due to longer sequence lengths. We set $\lambda=0.90$ for the T-REx experiments in §5 and $\lambda=0.99$ for the C4 experiments in §6. We also find that these values work well empirically, selecting from {0.50, 0.90, 0.99, 0.999}. Determining the optimal $\lambda$ for a given task (i.e. the number of task-specific components to downweight) is an interesting direction for future investigation.
We also note that this task-specific approach defines semantic ”dimensions” that should be downweighted (in this case, downweighting examples that support the overall task structure rather than an individual fact). When there is no a priori known task, the non-task-specific Hessian can be used, at a slight cost to performance (Experiment 5 in Table 1). More generally, the determination of semantic dimensions that should be downweighted is somewhat subjective; the eval data itself is not necessarily required for Hessian approximation, but rather any set of examples to define what semantics are less relevant.
A.2 Model Details
As described in §4.1, we pretrain a 154M-, 1B-, and 8B-parameter decoder-only language model on two epochs of English C4 (Raffel et al., 2020). All of our models are implemented using T5X (Roberts et al., 2022). For all model sizes, we use the same SentencePiece tokenizer (Kudo & Richardson, 2018) trained on C4 data with vocabulary size 32K. All model sizes are trained on the same shuffle of the pretraining data. We pretrain with batch size 1024 and sequence length 2048 for two epochs (187K steps). The 154M, 1B, and 8B models reach eval losses (log-perplexities) of $2.34$, $1.99$, and $1.77$ respectively. Specific hyperparameters are in Table A.2.